OpenStudy (anonymous):

Can anyone help me understand the coversion foctor?...for insistance...just an example...315 degrees *(pi/180)=7pi/4.

OpenStudy (kaiz122):

180 degrees = pi

OpenStudy (anonymous):

Yes, I get that one.

OpenStudy (kaiz122):

\[315^{o}* \frac{\pi}{180}=\frac{\pi}{4}\] we are converting \[315^{o}\] to rad, \[\frac{\pi}{180}\] means in a pi there is 180 degrees

OpenStudy (anonymous):

OK I get we are coverting. The soltion shuold be 7pi/4. Also, (please help me) understand why the bottom number is 4.

OpenStudy (kaiz122):

by simply dividing 350 by 180. you'll get 7/4

OpenStudy (kaiz122):